OpenStudy (anonymous):

Solve 3x - 5y = 15 solve for y

OpenStudy (anonymous):

3x - 5y = 15 -3x -3x (subtract 3x from both sides) -5y = 15 - 3x /-5 /-5 (divide both sides by -5) y = -3 + (3/5)y

OpenStudy (saifoo.khan):

3x - 15 = 5y y = 3x - 15 ------- 3

OpenStudy (anonymous):

3x-5y=15-------(i) or,3x=15+5y or,x=15+5y/3---------(ii) Now, keeping value of x in the equation----(i) 3x-5y=15 or,3*15+5y/3-5y=15 or,45+15y/3=15+5y or,45+15y=45+15y or,0=0